OpenStudy (anonymous):

Find the number a such that the line x = a divides the region bounded by the curves x = y2 − 1 and the y-axis into 2 regions with equal area. Give your answer correct to 3 decimal places.

OpenStudy (solomonzelman):

At first the full are that we are talking about: The boundaries are: \(\large\color{black}{ \displaystyle x=y^2-1 }\) \(\large\color{black}{ \displaystyle x=0 }\) ------------------------------------